OpenStudy (anonymous):

Estimate 1/1002 using differentials (or equivalently, a linear approximation). how do i start it if this is not a sqrt?

OpenStudy (anonymous):

how do you find the f(x)?

OpenStudy (phi):

how about f(x)= 1/x use a=1000 because it is close to 1002 but easy to find 1/1000 = 0.001